OpenStudy (anonymous):

Use the following equation: y = x2 The directrix is?

OpenStudy (anonymous):

@study100

OpenStudy (anonymous):

y= a(x-h)^2 +k directrix y= k- (1/4a) k=0 a=1 y= 0-(1/4) seems like directrix is y= -1/4 Sources and extra: this equation will set you for life with basic conics: http://www.jfkcougars.org/ourpages/auto/2013/2/5/43231143/Conics-Formula-Sheet.pdf

OpenStudy (anonymous):

So is it C?

OpenStudy (anonymous):

@study100

OpenStudy (anonymous):

I think so, since I haven't seen the answer choices. Got a pic?